5. Each firm in a competitive market has a cost function TC() 102q+ and there are n = 20 firms in the short run. The short-run equilibrium price is p-5. What happens in the long-run equilibrium as compared to the short-run equilibrium (a) Does the equilibrium price increase or decrease? Explain (b) Does the number of firms increase or decrease? Explain (c) Does the output of a single firm increase or decrease? Explain (d) Does the total market output increase or decrease? Explain.
Answer #1

From total cost, we have MC = 2 + 3q^2 and price is P = 5. Hence in the short run production is 5 = 2 + 3q^2 or q = 1 unit. Market quantity is 20 units. Long run price is P = MC = AC or 10/q + 2 + q^2 = 2 + 3q^2. This gives 10/q - 2q^2 or q = 1.70. Price is 10.77. Since short run price is less and long run price is higher, it appears that in the long run firms would leave the market since in the short run they are earning a loss. Number of firms will reduce and with the total quantity will also be.

a) Increase to 10.77

b) Decrease

c) Increase to 1.70

d) Decrease as supply curve will shift left.